OpenStudy (saylilbaby):

which cosine function has maximum of 4, a minimum of -4, and a period of 2pie/3?

OpenStudy (saylilbaby):

@dan815 @hartnn @idku

OpenStudy (boldjon):

period of cos x is 2π period of cos (ax) is 2π/a because cos (a(x + 2π/a) ) = cos (ax + 2π) = cos (ax) if we need period 2π/3 the function must be cos (3x) furthermore, as - 1 ≤ cos x ≤ 1 it will be - 2 ≤ 2 cos x ≤ 2 putting together the two things, the answer is c) y = 2 cos (3θ)

OpenStudy (boldjon):

de nada (¬‿¬)

OpenStudy (mathstudent55):